Stella is 4 years older than Maria. In 12 years, Stella will be twice Maria's present age. How old is Maria now?
Answer by ikleyn(52799)

.
Stella is 4 years older than Maria. In 12 years, Stella will be twice Maria's present age.
How old is Maria now?
~~~~~~~~~~~~~~~~~~~~

Let x be the Maria's present age.

Then the Stell's age is  (x+4), according to the problem's description.


In 12 years, Stella's age will be  ((x+4)+12) = (x+16) years.


According to the problem's description, we can write this equation


    x + 16 = 2x.


It implies


    16 = 2x - x

     x = 16.


ANSWER.  Maria is 16 years old.  It is her present age.

Solved.

-----------------

There is a bunch of lessons on age word problems
    - Age problems and their solutions
    - HOW TO algebreze and to solve age problems?
    - A fresh formulation of a traditional age problem
    - Really intricate age word problems
    - Selected age word problems from the archive
    - Age problems for mental solution
    - Age problem for three participants
    - Miscellaneous age problems
in this site.

Read them and become an expert in solving age problems.

Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Age word problems".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.